# Infrared-Assisted Single-Stage Framework for Joint Restoration and Fusion of Visible and Infrared Images under Hazy Conditions

### v1

@@ Abstract
Infrared and visible (IR-VIS) image fusion has gained significant attention for its broad application value. However, existing methods often neglect the complementary role of infrared image in restoring visible image features under hazy conditions. To address this, we propose a joint learning framework that utilizes infrared image for the restoration and fusion of hazy IR-VIS images. To mitigate the adverse effects of feature diversity between IR-VIS images, we introduce a prompt generation mechanism that regulates modality-specific feature incompatibility. This creates a prompt selection matrix from non-shared image information, followed by prompt embeddings generated from a prompt pool. These embeddings help generate candidate features for dehazing. We further design an infrared-assisted feature restoration mechanism that selects candidate features based on haze density, enabling simultaneous restoration and fusion within a single-stage framework. To enhance fusion quality, we construct a multi-stage prompt embedding fusion module that leverages feature supplementation from the prompt generation module. Our method effectively fuses IR-VIS images while removing haze, yielding clear, haze-free fusion results. In contrast to two-stage methods that dehaze and then fuse, our approach enables collaborative training in a single-stage framework, making the model relatively lightweight and suitable for practical deployment. Experimental results validate its effectiveness and demonstrate advantages over existing methods.
1 Introduction
@@
In recent years, significant progress has been made in IR-VIS image fusion; however, existing methods generally assume that the input visible images are of good visual quality. In hazy conditions, visible images are affected by haze, resulting in unclear imagery, which makes it difficult for these methods to generate clear, haze-free fusion results. Traditional approaches typically address this issue by first applying a dehazing algorithm to the hazy image and then fusing the dehazed image with the infrared image, as shown in Figure 1(a). Although this two-stage strategy is feasible, it fails to integrate the dehazing and fusion tasks into a unified framework for joint training, making it challenging to balance the relationship between the two tasks. While dehazed images may show good dehazing performance, they are not always optimal for subsequent fusion tasks. Additionally, the two-stage process of dehazing followed by fusion involves different methodologies, reducing the model’s compactness.
@@
In response to the challenge of IR-VIS image fusion and restoration under hazy conditions, we propose an infrared-assisted joint learning framework, as shown in Figure 1(b). To mitigate the impact of discrepancies between IR-VIS images on hazy image feature restoration, we design a prompt generation mechanism. It leverages non-shared information from input images to create a prompt selection matrix that selects and generates prompt embeddings from a prompt pool. These embeddings act as candidate features to aid in the recovery of hazy image features. For effective restoration of haze-affected features, we construct an infrared-assisted feature restoration module. It guides the selection of candidate features based on haze density to restore visible image features impacted by haze, enabling the joint processing of restoration and fusion within a single-stage framework. In this process, our focus shifts from solely enhancing the restoration of hazy visible images to emphasizing how restored features can further improve the quality of the fusion results.
@@
2 Related Work
2.1 Typical Fusion Methods
In IR-VIS image fusion, traditional methods based on multi-scale transforms and sparse representations [11, 14, 22, 46, 12] remain relevant. However, deep learning-based techniques have become mainstream. These methods can be categorized into three types: Convolutional Neural Network (CNN)-based methods [16, 23, 15], hybrid CNN-Transformer methods [17, 3, 33, 5], and Generative Adversarial Network (GAN)-based methods [25, 35, 26, 39, 21, 18, 47]. CNN-based methods extract features from input images and perform fusion using specialized modules, enhancing image details and contrast. However, CNNs are limited in modeling long-range dependencies, which impacts fusion quality in complex scenes. In contrast, Transformers excel at capturing long-range dependencies but struggle with local details and edges. Hybrid methods, such as AFT [3], YDTR [33], and HitFusion [5], combine CNN with Transformer to model local and global information, improving fusion performance.
Figure 2: Overall framework of the proposed method. The input IR and hazy VIS image pair $\left\{{{\bm{I}}_{ir}},{{\bm{I}}_{vi}}\right\}$ is processed by the PGM to obtain features $\left\{{{\bm{F}}_{ir}},{{\bm{F}}_{vi}}\right\}$ and a prompt ${{\bm{\hat{P}}}_{ir}}$ for ${{\bm{F}}_{ir}}$. Through the PEB, the prompt embedding ${{\bm{\hat{P}}}_{ir}}$ is used to refine the IR feature ${{\bm{F}}_{ir}}$, reducing redundant information and generating the refined IR feature ${{\hat{\bm{F}}_{ir}}}$. The HDE module [9] estimates the haze density in the VIS features to dynamically adjust the proportion of injected IR information, preventing excessive IR injection. The Transformer block removes degradation from the input features to obtain haze-free features. In the MsPE-FM, the haze-free VIS features and IR features are combined and passed to the Fusion Block for feature fusion. The PGM and PEB further are used to enhance the IR-VIS complementary information, reconstructing the final fused image.
@@ 2.1 Typical Fusion Methods
2.2 Methods Under Complex Imaging Conditions
@@
3 Methodology
3.1 Overview
As shown in Figure 2, our method comprises three core modules: the Infrared-Assisted Feature Restoration Module (IA-FRM), the Prompt Generation Module (PGM), and the Multi-stage Prompt Embedding Fusion Module (MsPE-FM). IA-FRM leverages infrared image features to assist in restoring lost information in heavily hazy regions of visible images, making it easier to restore these hazy areas. PGM generates a set of prompts to overcome the limitations of infrared images when assisting in the restoration of features in these dense hazy regions. Using the restored visible image features and prompts from PGM, MsPE-FM performs the fusion of IR-VIS image features, reconstructing a haze-free fused result.
3.2 Prompt generation
Infrared imaging sensors maintain performance in hazy conditions, allowing them to penetrate heavy haze. In a rigorously registered pair of IR-VIS images $\bm{I}_{ir},\bm{I}_{vi}$ provided to the model, we assume that only the visible image $\bm{I}_{vi}$ contains haze, while the infrared image $\bm{I}_{ir}$ is unaffected by haze. The core challenge of this work is to effectively utilize the infrared image $\bm{I}_{ir}$ to restore the visible image $\bm{I}_{vi}$ and then fuse them. However, the significant modality differences between IR-VIS images make it difficult to rely solely on the infrared image $\bm{I}_{ir}$ to recover the details lost in the hazy visible image. To overcome these challenges, this paper proposes the PGM, which generates a prompt embedding to create compensatory features that address the limitations of infrared features.
As shown in Figure 2, in the PGM, we first utilize an encoder constructed with Restormer [43] to perform feature encoding on the input registered IR-VIS images $\{{\bm{I}}_{ir},{\bm{I}}_{vi}\}$. As depicted in Figure 2(a), Restormer consists of a self-attention layer and a feed-forward network layer. The features output by the Restormer encoder are denoted as ${{\bm{F}}_{vi}}\in\mathbb{R}^{C\times H\times W}$ and ${{\bm{F}}_{ir}}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ represent the number of channels, height, and width of the features, respectively. Additionally, in this module, the features of the hazy visible image ${\bm{I}}_{vi}$ and the infrared image ${\bm{I}}_{ir}$ are processed by:
@@ 3.2 Prompt generation
3.3 Feature restoration assisted by infrared image
To effectively utilize the information provided by the infrared image $\bm{I}_{ir}$ for restoring features in haze regions, we design the IA-FRM. As shown in Figure 2, when restoring features in hazy images, regions with higher haze density should receive more focus. Therefore, it is essential to estimate the haze density in the input images. To achieve this, we adopt the method from [9] to estimate the haze density in the input image features, as illustrated in Figure 2(c). Specifically, the dark channel information of ${{\bm{F}}_{vi}}$ is obtained through the Dark Channel Generation (DCG) operation. The resulting dark channel information, along with the feature ${\tilde{\bm{F}}}_{vi}$, is then fed into the Atmospheric Light Estimation (ALE) module to predict the atmospheric light, denoted as $\bm{A}$. Here, ${\tilde{\bm{F}}}_{vi}$ is obtained by applying Channel Average Pooling (CAP) to ${{\bm{F}}_{vi}}$, followed by channel replication. Based on $\bm{A}$ and ${\tilde{\bm{F}}}_{vi}$, we can obtain the initial transmission map for each pixel location through the Get Haze Density (GHD) block, as shown in Eq. (3):
@@ 3.3 Feature restoration assisted by infrared image
3.4 Multi-stage Prompt Embedding Fusion
@@
As shown in Figure 2, in the MsPE-FM, the restored feature $\bm{\hat{F}}_{vi}$ and the initial feature $\bm{F}_{ir}$ are input into the PGM to obtain the prompt embedding $\bm{\hat{P}}_{vi}^{1}$ at the first stage of the fusion process. After concatenating $\bm{\hat{F}}_{vi}$ and $\bm{F}_{ir}$, the result is passed through the Restormer [43]. This output, along with $\bm{\hat{P}}_{vi}^{1}$, is then input into the PEB to obtain the fusion result for the next stage. In the second stage, $\bm{\hat{P}}_{vi}^{1}$ is first passed through a $1\times 1$ Conv layer to adjust the number of channels, resulting in the adjusted prompt embedding $\bm{\hat{P}}_{vi}^{2}$, which adapts to the changes in feature channels during the second-stage feature extraction. In this fusion process, five fusion blocks, each consisting of prompt embeddings, a Restormer, and a PEB, are used to achieve the fusion of IR-VIS features. Within these fusion blocks, two residual connections are employed to prevent information loss. Finally, the fused features pass through a $1\times 1$ Conv layer to reconstruct the fused result $\bm{I}_{f}$.
@@
4 Experiments
4.1 Experiment Settings
Datasets and Training Details. In this work, we utilize 1,083 IR-VIS image pairs from the MSRS dataset [31] as the training set. For the test set, we select 361 pairs from MSRS (ensuring no overlap with the training set), 50 pairs from RoadScene [36], and 100 pairs from M3FD [21]. To generate hazy image pairs, we apply the atmospheric scattering model [27] to introduce haze into the visible images in both the training and test sets. All experiments are conducted using the PyTorch framework on a single 24GB NVIDIA GeForce RTX 4090 GPU. During training, images are randomly cropped to $256\times 256$ patches, with data augmentation techniques such as horizontal and vertical flipping applied. The model is trained for a total of 300 epochs, using a batch size of 6 and the AdamW optimizer [10]. The initial learning rate is set to $2\times 10^{-4}$ and is gradually reduced to $2\times 10^{-6}$ following a cosine annealing schedule.
Figure 3: Visual comparison of fusion results from different methods. In the fusion results produced by the comparison methods, the first row of each image pair (except for the last column) shows the result after dehazing with DIACMP, followed by fusion. The second row shows the result after dehazing with Dehazeformer, followed by fusion. The last column presents the fusion results of Text-IF and our method.
Figure 4: Visualization of objective evaluation results. The comparison is divided into three groups: the first group shows the results of using DIACMP for dehazing followed by fusion on MSRS, M3FD, and RoadScene; the second group shows the results of using Dehazformer for dehazing followed by fusion on the same datasets; the third group presents the objective evaluation results of fusion images obtained using different dehazing methods, assessed with haze evaluation metrics.
Figure 5: Effect of different modules on fusion performance.
@@ 4.1 Experiment Settings
Metrics. To objectively evaluate the fusion performance of different methods, we adopt five commonly used image quality assessment metrics: Mutual Information ($Q_{MI}$) [29], Gradient-based Fusion Performance ($Q_{AB/F}$) [37], Chen-Varshney Metric ($Q_{CV}$) [4], Sum of Correlation of Differences ($Q_{SCD}$) [1], and Visual Information Fidelity ($Q_{VIF}$) [8]. These metrics assess the quality of the fusion results, using clear source images (without haze) as reference images when necessary. Additionally, to evaluate the perceptual quality of the dehazing effects within the fusion results, we employ the Perceptual Index ($Q_{PI}$) [2], Natural Image Quality Evaluator ($Q_{NIQE}$) [28], and Spatial Frequency ($Q_{SF}$) [7]. According to the evaluation criteria, lower $Q_{CV}$, $Q_{PI}$, and $Q_{NIQE}$ values indicate better fusion performance, while higher values for the remaining metrics signify improved quality.
4.2 Comparison with State-of-the-art Methods

### v2

@@ Abstract
Infrared and visible (IR-VIS) image fusion has gained significant attention for its broad application value. However, existing methods often neglect the complementary role of infrared image in restoring visible image features under hazy conditions. To address this, we propose a joint learning framework that utilizes infrared image for the restoration and fusion of hazy IR-VIS images. To mitigate the adverse effects of feature diversity between IR-VIS images, we introduce a prompt generation mechanism that regulates modality-specific feature incompatibility. This creates a prompt selection matrix from non-shared image information, followed by prompt embeddings generated from a prompt pool. These embeddings help generate candidate features for dehazing. We further design an infrared-assisted feature restoration mechanism that selects candidate features based on haze density, enabling simultaneous restoration and fusion within a single-stage framework. To enhance fusion quality, we construct a multi-stage prompt embedding fusion module that leverages feature supplementation from the prompt generation module. Our method effectively fuses IR-VIS images while removing haze, yielding clear, haze-free fusion results. In contrast to two-stage methods that dehaze and then fuse, our approach enables collaborative training in a single-stage framework, making the model relatively lightweight and suitable for practical deployment. Experimental results validate its effectiveness and demonstrate advantages over existing methods. The source code of the paper is available at https://github.com/fangjiaqi0909/IASSF.
Index Terms:
IR-VIS Image Fusion, Haze Removal, Joint Restoration and Fusion.
I Introduction
@@
In recent years, significant progress has been made in IR-VIS image fusion[1, 2, 3, 4, 5]; however, existing methods generally assume that the input visible images are of good visual quality. In hazy conditions, visible images are affected by haze, resulting in unclear imagery, which makes it difficult for these methods to generate clear, haze-free fusion results. Traditional approaches typically address this issue by first applying a dehazing algorithm to the hazy image and then fusing the dehazed image with the infrared image, as shown in Fig. 1(a). Although this two-stage strategy is feasible, it fails to integrate the dehazing and fusion tasks into a unified framework for joint training, making it challenging to balance the relationship between the two tasks. While dehazed images may show good dehazing performance, they are not always optimal for subsequent fusion tasks. Additionally, the two-stage process of dehazing followed by fusion involves different methodologies, reducing the model’s compactness.
@@
In response to the challenge of IR-VIS image fusion and restoration under hazy conditions, we propose an infrared-assisted joint learning framework, as shown in Fig. 1(b). To mitigate the impact of discrepancies between IR-VIS images on hazy image feature restoration, we design a prompt generation mechanism. It leverages non-shared information from input images to create a prompt selection matrix that selects and generates prompt embeddings from a prompt pool. These embeddings act as candidate features to aid in the recovery of hazy image features. For effective restoration of haze-affected features, we construct an infrared-assisted feature restoration module. It guides the selection of candidate features based on haze density to restore visible image features impacted by haze, enabling the joint processing of restoration and fusion within a single-stage framework. In this process, our focus shifts from solely enhancing the restoration of hazy visible images to emphasizing how restored features can further improve the quality of the fusion results.
@@
Figure 2: Overall framework of the proposed method. The input IR and hazy VIS image pair $\left\{{{\bm{I}}_{ir}},{{\bm{I}}_{vi}}\right\}$ is processed by the PGM to obtain features $\left\{{{\bm{F}}_{ir}},{{\bm{F}}_{vi}}\right\}$ and a prompt ${{\bm{\hat{P}}}_{ir}}$ for ${{\bm{F}}_{ir}}$. Through the PEB, the prompt embedding ${{\bm{\hat{P}}}_{ir}}$ is used to refine the IR feature ${{\bm{F}}_{ir}}$, reducing redundant information and generating the refined IR feature ${{\hat{\bm{F}}_{ir}}}$. The haze density esitimation (HDE) module [8] estimates the haze density in the VIS features to dynamically adjust the proportion of injected IR information, preventing excessive IR injection. The Transformer block removes degradation from the input features to obtain haze-free features. In the MsPE-FM, the haze-free VIS features and IR features are combined and passed to the Fusion Block for feature fusion. The PGM and PEB further are used to enhance the IR-VIS complementary information, reconstructing the final fused image.
II Related Work
II-A Typical Fusion Methods
In IR-VIS image fusion, traditional methods based on multi-scale transforms and sparse representations [9, 10, 11, 12, 13] remain relevant. However, deep learning-based techniques have become mainstream. These methods can be categorized into three types: Convolutional Neural Network (CNN)-based methods [14, 15, 16, 17], hybrid CNN-Transformer methods [18, 19, 20, 21], and Generative Adversarial Network (GAN)-based methods [22, 23, 24, 25, 26, 27, 4]. CNN-based methods extract features from input images and perform fusion using specialized modules, enhancing image details and contrast. However, CNNs are limited in modeling long-range dependencies, which impacts fusion quality in complex scenes. In contrast, Transformers excel at capturing long-range dependencies but struggle with local details and edges. Hybrid methods, such as AFT [19], YDTR [20], and HitFusion [21], combine CNN with Transformer to model local and global information, improving fusion performance.
@@ II-A Typical Fusion Methods
II-B Methods Under Complex Imaging Conditions
@@
III The Proposed Method
III-A Overview
As shown in Fig. 2, our method comprises three core modules: the Infrared-Assisted Feature Restoration Module (IA-FRM), the Prompt Generation Module (PGM), and the Multi-stage Prompt Embedding Fusion Module (MsPE-FM). IA-FRM leverages infrared image features to assist in restoring lost information in heavily hazy regions of visible images, making it easier to restore these hazy areas. PGM generates a set of prompts to overcome the limitations of infrared images when assisting in the restoration of features in these dense hazy regions. Using the restored visible image features and prompts from PGM, MsPE-FM performs the fusion of IR-VIS image features, reconstructing a haze-free fused result.
III-B Prompt Generation
Infrared imaging sensors maintain performance in hazy conditions, allowing them to penetrate heavy haze. In a rigorously registered pair of IR-VIS images $\{\bm{I}_{ir},\bm{I}_{vi}\}$ provided to the model, we assume that only the visible image $\bm{I}_{vi}$ contains haze, while the infrared image $\bm{I}_{ir}$ is unaffected by haze. The core challenge of this work is to effectively utilize the infrared image $\bm{I}_{ir}$ to restore the visible image $\bm{I}_{vi}$ and then fuse them. However, the significant modality differences between IR-VIS images make it difficult to rely solely on the infrared image $\bm{I}_{ir}$ to recover the details lost in the hazy visible image. To overcome these challenges, this paper proposes the PGM, which generates a prompt embedding to create compensatory features that address the limitations of infrared features.
As shown in Fig. 2, in the PGM, we first utilize an encoder constructed with Restormer [36] to perform feature encoding on the input registered IR-VIS images $\{{\bm{I}}_{ir},{\bm{I}}_{vi}\}$. As depicted in Fig. 2(a), Restormer consists of a self-attention layer and a feed-forward network layer. The features output by the Restormer encoder are denoted as ${{\bm{F}}_{vi}}\in\mathbb{R}^{C\times H\times W}$ and ${{\bm{F}}_{ir}}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ represent the number of channels, height, and width of the features, respectively. Additionally, in this module, the features of the hazy visible image ${\bm{I}}_{vi}$ and the infrared image ${\bm{I}}_{ir}$ are processed by:
@@ III-B Prompt Generation
III-C Feature Restoration Assisted by Infrared Image
To effectively utilize the information provided by the infrared image $\bm{I}_{ir}$ for restoring features in haze regions, we design the IA-FRM. As shown in Fig. 2, when restoring features in hazy images, regions with higher haze density should receive more focus. Therefore, it is essential to estimate the haze density in the input images. To achieve this, we adopt the method from [8] to estimate the haze density in the input image features, as illustrated in Fig. 2(c). Specifically, the dark channel information of ${{\bm{F}}_{vi}}$ is obtained through the Dark Channel Generation (DCG) operation. The resulting dark channel information, along with the feature ${\tilde{\bm{F}}}_{vi}$, is then fed into the Atmospheric Light Estimation (ALE) module to predict the atmospheric light, denoted as $\bm{A}$. Here, ${\tilde{\bm{F}}}_{vi}$ is obtained by applying Channel Average Pooling (CAP) to ${{\bm{F}}_{vi}}$, followed by channel replication. Based on $\bm{A}$ and ${\tilde{\bm{F}}}_{vi}$, we can obtain the initial transmission map for each pixel location through the Get Haze Density (GHD) block, as shown in Eq. (3):
@@ III-C Feature Restoration Assisted by Infrared Image
III-D Multi-stage Prompt Embedding Fusion
@@
As shown in Fig. 2, in the MsPE-FM, the restored feature $\bm{\hat{F}}_{vi}$ and the initial feature $\bm{F}_{ir}$ are input into the PGM to obtain the prompt embedding $\bm{\hat{P}}_{vi}^{1}$ at the first stage of the fusion process. After concatenating $\bm{\hat{F}}_{vi}$ and $\bm{F}_{ir}$, the result is passed through the Restormer [36]. This output, along with $\bm{\hat{P}}_{vi}^{1}$, is then input into the PEB to obtain the fusion result for the next stage. In the second stage, $\bm{\hat{P}}_{vi}^{1}$ is first passed through a $1\times 1$ Conv layer to adjust the number of channels, resulting in the adjusted prompt embedding $\bm{\hat{P}}_{vi}^{2}$, which adapts to the changes in feature channels during the second-stage feature extraction. In this fusion process, five fusion blocks, each consisting of prompt embeddings, a Restormer, and a PEB, are used to achieve the fusion of IR-VIS features. Within these fusion blocks, two residual connections are employed to prevent information loss. Finally, the fused features pass through a $1\times 1$ Conv layer to reconstruct the fused result $\bm{I}_{f}$.
@@
IV Experiments
IV-A Experimental Configurations
Dataset. In this work, we utilize 1,083 IR-VIS image pairs from the MSRS dataset [28] as the training set. This dataset includes a wide variety of scenes, such as vehicles, pedestrians, houses, and streets, offering a rich and diverse set of visual data for training purposes. For testing, we use 361 image pairs from the MSRS dataset for both qualitative and quantitative comparative experiments, ensuring no overlap with the training set. Additionally, we assess the effectiveness and generalization capability of our method on 100 image pairs from the M3FD dataset [26] and 50 image pairs from the RoadScene dataset [38]. The M3FD dataset contains IR-VIS image pairs from scenes such as university campuses, vacation spots, and urban main roads, while the RoadScene dataset includes IR-VIS image pairs selected from the representative scenes including pedestrians, vehicles, roads, and buildings. To generate hazy image pairs, we apply the atmospheric scattering model [8] to introduce haze into the visible images in both the training and test sets.
@@ IV-A Experimental Configurations
Metrics. To objectively evaluate the fusion performance of different methods, we adopt five commonly used image quality assessment metrics: Mutual Information ($Q_{MI}$) [39], Gradient-based Fusion Performance ($Q_{AB/F}$) [40], Chen-Varshney Metric ($Q_{CV}$) [41], Sum of Correlation of Differences ($Q_{SCD}$) [42], and Visual Information Fidelity ($Q_{VIF}$) [43]. These metrics are used to assess the quality of the fusion results, with clear source images (without haze) as reference images when necessary. Additionally, to evaluate the perceptual quality of the dehazing effects within the fusion results, we employ the Perceptual Index ($Q_{PI}$) [44], Natural Image Quality Evaluator ($Q_{NIQE}$) [45], and Spatial Frequency ($Q_{SF}$) [46]. According to the evaluation criteria, lower $Q_{CV}$, $Q_{PI}$, and $Q_{NIQE}$ values indicate better fusion performance, while higher values for the remaining metrics signify improved quality.
IV-B Implementation Details
All experiments are conducted using the PyTorch framework on a single 24GB NVIDIA GeForce RTX 4090 GPU. During training, images are randomly cropped to $256\times 256$ patches, with data augmentation techniques such as horizontal and vertical flipping applied. The model is trained for a total of 300 epochs, using a batch size of 6 and the AdamW optimizer [47]. The initial learning rate is set to $2\times 10^{-4}$ and is gradually reduced to $2\times 10^{-6}$ following a cosine annealing schedule.
IV-C Comparison with State-of-the-art Methods